OpenStudy (curry):

find the parametrization for the circle with center at (-3,2) and with radius = 3

OpenStudy (dumbcow):

start with rectangular equation for circle \[(x+3)^{2} + (y-2)^{2} = 9\] then note that goal is to get it in this form: 9(sin^2 + cos^2) = 9 since the trig identity sin^2 + cos^2 = 1 \[x = 3\cos(t) -3\] \[y=3\sin(t)+2\] if you plug these values in you get: \[9\cos^{2} (t) + 9\sin^{2} (t) = 9\]

OpenStudy (dumbcow):

in general for any circle with center (h,k) \[x = r*\cos(t) +h\] \[y = r*\sin(t)+k\]

OpenStudy (curry):

oh sso x=3cos -3 and y=3sin + 2 is the ansswer ... i thoiught u had to have it equaling t thatr what i did wrong

OpenStudy (curry):

ty ssmart cow